1. If a company sets aside $1,000,000 now into a contingency fund, how much will the company have in 2 years, if it does not use any of the money and the account grows at a rate of 10% per year?


1
Expert's answer
2021-12-20T19:57:13-0500

F1 = 1,000,000 + 1,000,000(0.10) = 1,100,000

F2 = 1,100,000 + 1,100,000(0.10) = $1,210,000